Be sure to answer all parts.

Calculate the amount of water (in grams) that must be added to each of the following substances to produce the indicated solution.

(a) 7.70 g of urea (NH2)2CO in the preparation of a 20.1 percent by mass solution:

(b) 22.4 g of MgCl2 in the preparation of a 6.6 percent mass solution:
Answer #1

a) Mass percent = ( mass of urea / mass of solution ) x 100

20.1 / 100 = 7.70 g / mass of solution

0.201 = 7.70 / mass of solution

mass of solution = 7.70 / .201

= 38.31g

Mass of water = mass of solution - mass of urea

= 38.31 - 7.70 = 30.61 g

Mass of water added = 30.61 g

b) Mass percent = ( mass of MgCl2 / mass of solution ) x 100

6.6 = ( 22.4 / mass of solution ) x 100

6.6/100 = 22.4 / mass of solution

.066 = 22.4 / mass of solution

mass of solution = 22.4 / .066

= 339.39 g

mass of water = mass of solution - mass of MgCl2

= 339.39 - 22.4

= 316.99 g

mass of water added = 316.99 g